OpenStudy (h0pe):

If \(554_b\) is the base \(b\) representation of the square of the number whose base \(b\) representation is \(24_b,\) then find \(b\).

myininaya (myininaya):

wait so it kind of looks like it is saying this: \[554_b=(24_b) \cdot (24_b)\]

myininaya (myininaya):

and of course we are to find the b there

OpenStudy (h0pe):

I think that looks about right

myininaya (myininaya):

\[5 \cdot b^2+ 5 \cdot b+4=(2 \cdot b+4)(2 \cdot b +4)\]

myininaya (myininaya):

this looks like a quadratic equation

OpenStudy (h0pe):

so then \((2b+4)(2b+4)\) is \(4b^2+8b+8\) right?

myininaya (myininaya):

well not exactly 4(4)=16 and 2b(4)+4(2b)=16b

myininaya (myininaya):

check that middle term and last term you have

OpenStudy (h0pe):

They are 16s whoops \(4b2+16b+16\)

OpenStudy (h0pe):

\(4b^2\)

myininaya (myininaya):

right so you have \[5b^2+5b+4=4b^2+16b+16\]

myininaya (myininaya):

you need to some subtraction on both sides

OpenStudy (h0pe):

okay

OpenStudy (h0pe):

so \(b^2=11b+12\)

myininaya (myininaya):

\[\text{ or } b^2-11b-12=0\]

myininaya (myininaya):

which you can factor the left hand side

myininaya (myininaya):

or you can use the quadratic formula if you really want to but factoring the left hand expression is not too bad

myininaya (myininaya):

you will get one b that makes sense and the other b that makes no sense

OpenStudy (h0pe):

what do you mean by factoring? sorry it's been a while since I've used that. 0.0

myininaya (myininaya):

like to factor something like x^2-5x+6 you can look for two numbers that multiply to be 6 and add up to be -5 so -2(-3)=-6 and -2+(-3)=-5 so x^2-5x+6 can be written as (x-2)(x-3) and if you wanted to solve x^2-5x+6=0 you can write this first as (x-2)(x-3)=0 then set both factors equal to zero x-2=0 or x-3=0 so x=2 or x=3 you can find two numbers that multiply to be -12 and add up to be -11

OpenStudy (h0pe):

Right, I remember now

OpenStudy (h0pe):

For the case of \(b^2−11b−12=0\) it would be -12 and 1, wouldn't it be? So the equation would be \((b-12)(b+1)\) right?

myininaya (myininaya):

bingo so you have to solve (b-12)(b+1)=0 which means you have either b-12=0 or b+1=0 (or both <--both will not work in this case and I hope you will see why )

OpenStudy (h0pe):

The answer has to be \(b=12\), because there aren't negative bases.

myininaya (myininaya):

cool stuff :)

OpenStudy (h0pe):

Thanks :)